# A Formal Comparison Between Chain-of-Thought and Latent Thought

### v1

@@ D.1.2 Self-Consistency Enhances Reliability
Fix $x,\varepsilon,\delta$. Run $A$ independently $k$ times with fresh randomness and the same $(x,\varepsilon)$, obtaining estimates $\widehat{f}^{(1)},\dots,\widehat{f}^{(k)}$. Output the median
$\widehat{f}_{\mathrm{med}}\;\coloneqq\;\mathrm{median}\!\left\{\widehat{f}^{(1)},\dots,\widehat{f}^{(k)}\right\}.$ (65)
For each run $i\in\{1,\dots,k\}$, define the Bernoulli indicator
$Y_{i}\;\coloneqq\;\mathbf{1}\!\left[(1-\varepsilon)f(x)\ \leq\ \widehat{f}^{(i)}\ \leq\ (1+\varepsilon)f(x)\right]\in\{0,1\},$ (66)
so that $p\coloneqq\Pr[Y_{i}=1]\geq\tfrac{1}{2}+\gamma$ by assumption, and let $S\coloneqq\sum_{i=1}^{k}Y_{i}$, so $\mathbb{E}[S]=pk$. If at least half the runs are successful (i.e., $S\geq\lceil k/2\rceil$), then at least $\lceil k/2\rceil$ samples lie inside the interval $\mathcal{I}=[(1-\varepsilon)f(x),(1+\varepsilon)f(x)]$, hence the median must also lie in $\mathcal{I}$. Therefore
$\Pr\!\left[\widehat{f}_{\mathrm{med}}\notin\mathcal{I}\right]\;\leq\;\Pr\!\left[S<\frac{k}{2}\right].$ (67)
Since the $Y_{i}$ are independent in $[0,1]$, Hoeffding’s inequality gives
$\Pr\left[S\leq\frac{k}{2}\right]\;=\;\Pr\left[\frac{S}{k}-p\leq-\Bigl(p-\frac{1}{2}\Bigr)\right]\;\leq\;\exp\bigl(-2k\,(p-\tfrac{1}{2})^{2}\bigr)\;\leq\;\exp\bigl(-2k\,\gamma^{2}\bigr).$ (68)
Choose $k\;\geq\;\frac{1}{2\gamma^{2}}\,\ln\!\frac{1}{\delta},$ then $\Pr[\widehat{f}_{\mathrm{med}}\notin\mathcal{I}]\leq\delta$. We invoke $A$ independently $k=O\bigl(\gamma^{-2}\log(1/\delta)\bigr)$ times and take a median, Hence the total running time is polynomial in $|x|$, $1/\varepsilon$, and $\log(1/\delta)$. ∎
Lemma D.4.
For $m\geq 1$ and $m\leq\varepsilon\leq\sqrt{m}$, it holds that
$e^{\,\varepsilon-\frac{1}{2}}\;\leq\;\Bigl(1+\frac{\varepsilon}{m}\Bigr)^{m}\;\leq\;e^{\,\varepsilon}.$ (69)
Proof.
For u>1u>-1, the elementary bounds $\ln(1+u)\leq u$ and $\ln(1+u)\geq u-\tfrac{u^{2}}{2}$ hold. Substituting $u=\varepsilon/m$ and multiplying by $m$ gives
$m\ln\Bigl(1+\frac{\varepsilon}{m}\Bigr)\leq\varepsilon,\qquad m\ln\Bigl(1+\frac{\varepsilon}{m}\Bigr)\geq\varepsilon-\frac{\varepsilon^{2}}{2m}.$ (70)
This yields, respectively,
$(1+\varepsilon/m)^{m}\leq e^{\varepsilon},\qquad(1+\varepsilon/m)^{m}\geq\exp\bigl(\varepsilon-\varepsilon^{2}/(2m)\bigr).$ (71)
If $\varepsilon^{2}\leq m$, then $\varepsilon^{2}/(2m)\leq\tfrac{1}{2}$, thus we have $(1+\varepsilon/m)^{m}\geq e^{\varepsilon-1/2}.$
D.1.3 Self-Reducibility
Definition D.5 (Schnorr, 1976).
@@ D.1.3 Self-Reducibility
For self-reducible relations, the existence of an FPRAS and an FPAUS is equivalent.
Theorem D.6 (Jerrum et al., 1986).
@@
D.2 Proof for Proposition 4.6
The basic strategy follows the reduction from almost uniform generation to approximate counting (Jerrum et al., 1986).
Proof.
Fix target parameters $\varepsilon,\delta>0$ for the FPAUS. Let $m\in\mathsf{poly}(n)$ be such that for any $y\in R(x)$ we have $|y|\leq m(|x|)$. By Lemma D.3 and the assumption of CoT model $\pi$, for every position $i\in[m(|x|)]$, every input $x\in\Sigma^{*}$, and every prefix $y_{<i}$, it holds that
$\Pr_{\omega}\!\left[(1-\tfrac{1}{2m(|x|)})\,p(y_{i}\mid x,y_{<i})\;\leq\;\pi_{\omega}(y_{i}\mid x,y_{<i})\;\leq\;(1+\tfrac{1}{m(|x|)})\,p(y_{i}\mid x,y_{<i})\right]\;\geq\;1-\frac{\delta}{3m(|x|)}\,$, (76)
where $\pi_{\omega}$ denotes the distribution obtained by taking the median over multiple trials, and $\omega$ represents the underlying randomness of these trials.
Hence the autoregressive sampling of $y_{1},\ldots,y_{m(|x|)}$ succeeds at all steps with probability at least
$1-m(|x|)\cdot\tfrac{\delta}{3m(|x|)}\;=\;1-\tfrac{\delta}{3}\,$, (77)
by the union bound. Condition on this “good” event. Let $p(y\mid x)=1/|R(x)|$ denote the uniform target on $R(x)$. Multiplying the stepwise bounds and using Lemma D.4, namely $\bigl(1+\tfrac{1}{2m}\bigr)^{m}\leq\mathrm{e}^{1/2}$, we obtain
$\mathrm{e}^{-1/2}\;\leq\;\frac{\pi_{\omega}(y\mid x)}{p(y\mid x)}\;\leq\;\mathrm{e}^{1/2}\qquad\text{for all }y\in R(x).$ (78)
To obtain exact uniformity, we apply rejection sampling. Run one preliminary draw $z\sim\pi_{\bm{\omega}}(\cdot\mid x)$ and set $\varphi_{0}\;\coloneq\;\mathrm{e}^{-1}\,\pi_{\bm{\omega}}(z\mid x),$ and define the acceptance probability, denoted by $a(y)$, as
$a(y)\;\coloneq\;\min\Bigl\{1,\,\frac{\varphi_{0}}{\pi_{\bm{\omega}}(y\mid x)}\Bigr\}.$ (79)
By equation 78 applied to both $z$ and any $y\in R(x)$,
$\pi_{\bm{\omega}}(z\mid x)\;\leq\;\mathrm{e}^{1/2}\,p(z\mid x)\;=\;\frac{\mathrm{e}^{1/2}}{|R(x)|}\quad\text{and}\quad\pi_{\bm{\omega}}(y\mid x)\;\geq\;\mathrm{e}^{-1/2}\,p(y\mid x)\;=\;\frac{\mathrm{e}^{-1/2}}{|R(x)|}.$ (80)
Therefore
$\varphi_{0}\;=\;\mathrm{e}^{-1}\pi_{\bm{\omega}}(z\mid x)\;\leq\;\frac{\mathrm{e}^{-1/2}}{|R(x)|}\;\leq\;\pi_{\bm{\omega}}(y\mid x)\qquad\text{for all }y\in R(x).$ (81)
Thus $a(y)$ is well-defined, and the output probabilities of the rejection sampling procedure satisfy
$\Pr[y]\;\propto\;\pi_{\bm{\omega}}(y\mid x)\,a(y)\;=\;\min\{\pi_{\bm{\omega}}(y\mid x),\,\varphi_{0}\}.$ (82)
which is constant in $y$. Hence, conditional on the good event, the output distribution is exactly uniform over $R(x)$. Still, on the good event, the acceptance probability in one trial is
$\Pr[\mathrm{accept}]\;=\;\sum_{y\in R(x)}\pi_{\bm{\omega}}(y\mid x)\,a(y)\;=\;|R(x)|\,\varphi_{0}\;\geq\;|R(x)|\cdot\mathrm{e}^{-1}\cdot\frac{\mathrm{e}^{-1/2}}{|R(x)|}\;=\;\mathrm{e}^{-3/2}.$ (83)
Hence, the per-trial rejection probability is at most $q\coloneq 1-\mathrm{e}^{-3/2}$. Repeating independently $k$ times, the probability of rejecting in all trials is at most $q^{k}$. To make this at most $\delta/3$, it suffices to take
$k\;\geq\;\log(3/\delta)/\log\bigl(1/q\bigr)\;=\;O(\log(1/\delta)).$ (84)
There are three failure modes: (i) stepwise approximation fails (probability $\leq\delta/3$); (ii) the preliminary draw does not satisfy equation 78 (already covered by (i), so no extra cost); (iii) all $k$ rejection-sampling trials reject ($\leq\delta/3$). By the union bound, the overall failure probability is at most $\delta$. On success, the output law is uniform, so its total variation distance to the uniform target is also at most $\delta$. The running time is polynomial in $|x|$ and $\log(1/\delta)$, so the procedure is an FPAUS. ∎
D.3 Proof for Theorem 4.7
First, we show the expressiveness of CoT with stochastic decoding, which can simulate a probabilistic Turing machine (PTM). For simplicity, we consider a two-tape PTM.
Definition D.7.
A probabilistic Turing machine $M=(Q,\Sigma,\Gamma,q_{0},q_{1},A,\delta_{1},\delta_{2})$ is defined as:
  • $Q$ is a finite set of states,
  • $\Sigma$ is the input/output alphabet,
Our result builds on Nowak et al. (2024), who extended that of Pérez et al. (2021) from simulating deterministic Turing machines with attention to incorporating probabilistic decoding, thereby capturing the probabilistic transitions of PTMs. Their proof, however, relies on hard attention, whereas our setting uses softmax attention. In addition, the definition of the output format of CoT models considered here differs from theirs. We therefore establish the corresponding statements in our setting.
Lemma D.8.
$\Pr\bigl[\mathsf{CoT}_{M}(x)=y\bigr]\;=\;\Pr\bigl[M(x)=y\bigr].$
Moreover, the number of reasoning steps of $\mathsf{CoT}_{M}$ is bounded by $O(T(|x|))$.
Proof.
We argue by induction on the number of machine steps and construct a CoT whose token sequence mirrors the PTM’s execution. Let $Q$ be the state set of $M$, $\Gamma$ the work-tape alphabet, and $\Sigma$ the input/output alphabet. At each step $i\geq 0$, the machine has a configuration summarized by
${\bm{r}}_{i}\;=\;\bigl(q^{(i)},\,v^{(i)},\,w^{(i)},\,m^{(i)}\bigr)\;\in\;Q\times\Gamma\times\bigl(\Sigma\cup\{\varepsilon\}\bigr)\times\{-1,0,1\},$
where $q^{(i)}$ is the current state, $v^{(i)}$ is the symbol written on the work tape (at the head cell) during step $i$, $w^{(i)}$ is the symbol written to the output tape (or $\varepsilon$ if no output), and $m^{(i)}\in\{-1,0,1\}$ is the head move. Let the output string be $y=(y_{1},\ldots,y_{m})\in\Sigma^{*}$, and let $r:[m]\to\mathbb{N}$ map each output index $i$ to the index of step $r(i)$ in which $M$ writes $y_{i}$ to the output tape.
The precision of CoT is $s(n)\;=\;C\cdot\bigl\lceil\log_{2}n\bigr\rceil,$ where $C\in\mathbb{N}$ is chosen sufficiently large so that $n^{C}$ exceeds the maximum polynomial step bound under consideration. For input $x=(x_{1},\ldots,x_{n})\in\Sigma^{n}$, the model produces the sequence
$\Bigl(x,\;{\bm{r}}_{1:r(1)},\;e,\;y_{1},\;e^{\prime},\;{\bm{r}}_{r(1)}^{\prime},\;{\bm{r}}_{r(1)+1:r(2)},\;e,\;y_{2},\;e^{\prime},\;{\bm{r}}_{r(2)}^{\prime},\;\ldots,\;{\bm{r}}_{r(m-1)+1:r(m)},\;e,\;y_{m},\;e^{\prime}\Bigr),$ (85)
where blocks of configuration tokens ${\bm{r}}_{i}$ simulate the PTM step-by-step and the triple $(e,y_{i},e^{\prime})$ deterministically realizes the $i$-th output when it occurs. The vocabulary of the model is
${\mathcal{V}}\;\coloneq\;Q\times\Gamma\times\bigl(\Sigma\cup\{\varepsilon\}\bigr)\times\{-1,0,1\}\times\{r,r^{\prime}\}\;\;\cup\;\;\bigl(\Sigma\cup\{\varepsilon\}\bigr)\;\;\cup\;\;\{e,e^{\prime}\}.$ (86)
Here, $\{r,r^{\prime}\}$ is a binary flag used to distinguish whether a configuration token carries the ordinary marker $r$ or the primed marker $r^{\prime}$.
Embeddings
Let $\bm{e}_{Q}:Q\to\{0,1\}^{|Q|}$, $\bm{e}_{\Gamma}:\Gamma\to\{0,1\}^{|\Gamma|}$, and $\bm{e}_{\Sigma}:\Sigma\to\{0,1\}^{|\Sigma|}$ denote one-hot encodings, and let ${\bm{p}}(\cdot)=\mathsf{sbin}_{s(n)}(\cdot)$ be the positional code. We write $\bm{0}$} for a zero block of appropriate size. The embedding of an input token $x_{i}$ is
$\bigl(\bm{e}_{Q}(q^{(0)}),\;\bm{e}_{\Gamma}(x_{i}),\;{\bm{0}}_{|\Sigma|},\;0,\;{\bm{p}}(i),\;{\bm{0}}\bigr).$ (87)
For a configuration token ${\bm{r}}_{i}=(q^{(i)},v^{(i)},w^{(i)},m^{(i)})$ at position $n+j$, the embedding is
$\bigl(\bm{e}_{Q}(q^{(i)}),\;\bm{e}_{\Gamma}(v^{(i)}),\;\bm{e}_{\Sigma}(w^{(i)}),\;m^{(i)},\;{\bm{p}}(j),\;{\bm{0}}\bigr).$ (88)
For delimiters $d\in\{e,e^{\prime}\}$ at position $n+j$, we use
[…]
Finally, for an output token $y_{i}\in\Sigma$ at position $n+j$, the embedding is
$\bigl({\bm{0}}_{|Q|},\;{\bm{0}}_{|\Gamma|},\;\bm{e}_{\Sigma}(y_{i}),\;0,\;{\bm{p}}(j),\;{\bm{0}},\;0,\;1,\;0,\;{\bm{0}}_{|{\mathcal{V}}|}\bigr).$ (90)
First layer
For each input position $i$, the model computes the head position of the Turing machine both at and immediately after the $i$-th step. This construction follows the approach of Pérez et al. (2021); Merrill & Sabharwal (2024); Nowak et al. (2024). The next head position is obtained by summing the head movement directions $m^{(j)}\in\{-1,0,1\}$ up to step $i$:
$c^{(i+1)}\;=\;m^{(0)}+m^{(1)}+\cdots+m^{(i)}.$ (91)
To realize this with single-head attention, we define the query, key, and value vectors at position $j$ as
$\mathbf{q}_{j}\;=\;(1),\quad\mathbf{k}_{j}\;=\;(1),\quad\mathbf{v}_{j}\;=\;(m^{(j)}).$ (92)
Since the attention scores are uniform across all positions, the attention layer computes the cumulative sum of head movements. A subsequent feed-forward layer then yields $c^{(i)}\;=\;c^{(i+1)}-m^{(i)}.$ Thus, at token ${\bm{r}}_{k}$ at position $n+i$, the representation is
$\Bigl(\,\bm{e}(q^{(k)}),\;\bm{e}(v^{(k)}),\;\bm{e}(w^{(k)}),\;m^{(k)},\;{\bm{p}}(i),\;c^{(k)}\;{\bm{0}}\Bigr).$ (93)
Second layer
For configuration ${\bm{r}}_{k}$, the model retrieves the symbol located under the head, denoted by $s^{(k)}$. To accomplish this, we define an attention mechanism that filters positions $j\leq k$ such that $c^{(j)}=c^{(k)}$, and selects the largest such index $j$. Formally, $s^{(k)}\;=\;v^{(j^{\star})},\quad\text{where}\quad j^{\star}=\max\{\,j\leq k:c^{(j)}=c^{(k)}\,\}.$ To implement this retrieval, we set the attention vectors at position $j$ as
$\mathbf{q}_{j}$ $=\Bigl({\mathsf{sbin}_{s(n)}(j)}^{\frown}{1_{s(n)}},\;T(n)\cdot j\Bigr),$ (94)
$\mathbf{k}_{j}$ $=2^{s(n)+1}\cdot\Bigl({\mathsf{sbin}_{s(n)}(j)}^{\frown}{-1_{s(n)}},\;T(n)\cdot j\Bigr),\quad\mathbf{v}_{j}=\Bigl(\bm{e}_{\Gamma}(v^{(j)}),\;{\bm{0}}\Bigr).$ (95)
Here, the construction guarantees that there are only $T(n)\cdot O(\log n)$ distinct scores. Although the resulting attention is not a perfect one-hot vector, the ordering of scores is preserved, so the maximum weight still corresponds to the correct index. Consequently, the model can reliably retrieve a value that approximates $\bm{e}_{\Gamma}(s^{(k)})$. The subsequent feed-forward layer computes the two transition functions $\delta_{1}$ and $\delta_{2}$ in parallel:
$z^{(k+1)}_{j}\;=\;\bm{e}\big(\delta_{j}\bigl(q^{(k)},\,s^{(k)}\bigr)\big)\;\coloneq\;\bigl(\bm{e}_{Q}(q^{(k+1)}_{j}),\,\bm{e}_{\Gamma}(v^{(k+1)}_{j}),\,\bm{e}_{\Sigma}(w^{(k+1)}_{j}),\,m^{(k+1)}_{j}\bigr),\quad j\in\{1,2\}.$ (96)
Finally, the overall output at token ${\bm{r}}_{k}$ at position $n+i$ is
$\Bigl(\,\bm{e}(q^{(k)}),\;\bm{e}(v^{(k)}),\;\bm{e}(w^{(k)}),\;m^{(k)},\;{\bm{p}}(i),\,c^{(k)},\,s^{(k)},\,z^{(k+1)}_{1},\,z^{(k+1)}_{2},\,{\bm{0}}\Bigr),$ (97)
where $z^{(k+1)}_{1}$ and $z^{(k+1)}_{2}$ represent the encodings of the two candidate transitions.
Third layer
This layer handles the delimiters. When the last input token is $e$, the model copies the most recently written output symbol $y_{i}$ from the immediately preceding position. When the last input token is $e^{\prime}$, the model copies the most recent configuration ${\textnormal{v}}_{i}$ from three positions earlier. We implement this mechanism using two attention heads:
$\mathbf{q}_{i,1}$ $\;=\;{\mathsf{sbin}_{s(n)}(i-1)}^{\frown}{1_{s(n)}},$ (98)
$\mathbf{q}_{i,2}$ $\;=\;{\mathsf{sbin}_{s(n)}(i-3)}^{\frown}{1_{s(n)}},$ (99)
$\mathbf{k}_{i,1}$ $\;=\;\mathbf{k}_{i,2}\;=\;2^{s(n)+1}\cdot{\mathsf{sbin}_{s(n)}(i)}^{\frown}{(-1_{s(n)})},$ (100)
$\mathbf{v}_{i,1}$ $\;=\;\mathbf{v}_{i,2}\;=\;\bigl(\bm{e}(q^{(i)}),\,\bm{e}(v^{(i)}),\,\bm{e}(w^{(i)}),\,m^{(i)},\,{\bm{0}}\bigr).$ (101)
The first head is configured to retrieve the symbol from one position earlier, while the second head retrieves the configuration from three positions earlier. The following feed-forward map, conditioned on the one-hot flags indicating whether the input token is one of $\{{\bm{r}}_{i},\,e,\,y_{i},\,e^{\prime}\}$, outputs $\bigl(\,*,\;{\bm{z}}\bigr),$ where ${\bm{z}}\in\{0,1\}^{|{\mathcal{V}}|}$ is defined casewise as
[…]
Here $\bm{e}(\cdot)$ denotes one-hot encodings aligned with the corresponding slots of $\mathcal{V}$}.
Output layer
The final output layer is $\mathbf{OUT}=\begin{bmatrix}{\bm{0}}&{\bm{I}}_{|{\mathcal{V}}|}\end{bmatrix}$. The decoding function generates tokens in two modes: when the current token is ${\bm{r}}_{i}$, it stochastically samples from the two transitions $z^{(k+1)}_{1}$ and $z^{(k+1)}_{2}$; otherwise, it deterministically outputs $y_{i}$ for $e$, $e^{\prime}$ for $y_{i}$, and ${\bm{r}}_{r(i)}$ for $e^{\prime}$.
Proof for Theorem 4.7.
Under the assumption $\mathsf{FPTAS}\subsetneq\mathsf{FPRAS}$ for self-reducible relations, let $R$ denote a self-reducible relation whose counting problem $|R(x)|$ admits an FPRAS but not an FPTAS.
Proof of the CoT Part
Define the extension counting function
$\operatorname{EXT}_{R}:\Sigma^{*}\times\Sigma^{*}\to\mathbb{N},\qquad\operatorname{EXT}_{R}(x,w)\;=\;\bigl|\{\,z\in\Sigma^{*}:\langle x,wz\rangle\in R\,\}\bigr|.$
First, we briefly show that if there exists an FPRAS for $|R(x)|$, then there also exists an FPRAS for $\operatorname{EXT}_{R}$, as shown by (Jerrum et al., 1986). Let $\mathcal{A}(x,\varepsilon,\delta)$ denote an FPRAS for the counting problem $|R(x)|$, i.e., for any $\varepsilon,\delta\in(0,1)$,
$\Pr\!\left[(1-\varepsilon)\,|R(x)|\;\leq\;\mathcal{A}(x,\varepsilon,\delta)\;\leq\;(1+\varepsilon)\,|R(x)|\right]\;\geq\;1-\delta.$ (103)
By self-reducibility, given $(x,w)$ we proceed as follows: while $|w|>\sigma(x)$, decompose $w=uv$ with $|u|=\sigma(x)$, set $x\leftarrow\psi(x,u)$ and $w\leftarrow v$. When the loop terminates, define $S:=\Sigma^{\sigma(x)-|w|}$. Then
$\operatorname{EXT}_{R}(x,w)\;=\;\sum_{u\in S}|R(\psi(x,wu))|.$
An $(\varepsilon,\delta)$-relative estimate of $\operatorname{EXT}_{R}(x,w)$ is obtained by replacing each summand with $\mathcal{A}\!\bigl(\psi(x,wu),\varepsilon,\delta/|S|\bigr)$ and summing. A union bound guarantees overall relative error within $(1\pm\varepsilon)$ with probability at least $1-\delta$. Let $\widehat{\operatorname{EXT}}_{R}(\cdot,\cdot)$ be these $(1\pm\varepsilon)$ estimates, and define
$\pi(y_{i}\mid x,y_{<i})\;:=\;\frac{\widehat{\operatorname{EXT}}_{R}(x,y_{<i}y_{i})}{\sum_{u\in\Sigma}\widehat{\operatorname{EXT}}_{R}(x,y_{<i}u)}\qquad(y_{i}\in\Sigma).$ (104)
The target (uniform) conditional distribution is

### v2

@@ C.3 Self-Reducibility and Complexity of Approximate Counting
We use the standard inequality $1+x\leq e^{x}$, which holds for all $x\in\mathbb{R}$. Substituting $x=\frac{\varepsilon}{2m}$, we have:
$\left(1+\frac{\varepsilon}{2m}\right)^{m}\leq\left(\exp\left(\frac{\varepsilon}{2m}\right)\right)^{m}=e^{\varepsilon/2}.$ (84)
For $0<\varepsilon\leq 1$, we show that $e^{\varepsilon/2}<1+\varepsilon$. Let $f(\varepsilon)=1+\varepsilon-e^{\varepsilon/2}$. Then $f(0)=0$ and $f^{\prime}(\varepsilon)=1-\frac{1}{2}e^{\varepsilon/2}$. Since $e^{\varepsilon/2}\leq e^{1/2}<2$ for all $\varepsilon\in[0,1]$, it follows that $f^{\prime}(\varepsilon)>0$. Thus, $f$ is strictly increasing on $[0,1]$, implying $f(\varepsilon)>f(0)=0$, or equivalently $e^{\varepsilon/2}<1+\varepsilon$. ∎
Proof for Proposition C.13.
Let $x\in\Sigma^{n}$ and $m=p(n)$. We express $N_{R}(x)$ as a product of $m$ ratios. Let $w=y_{1}y_{2}\dots y_{m}$ be a witness, and let $w^{(i)}$ denote its prefix of length $i$. We can write:
$N_{R}(x)=\mathrm{Ext}_{R}(x,\lambda)=\prod_{i=1}^{m}\frac{\mathrm{Ext}_{R}(x,w^{(i-1)})}{\mathrm{Ext}_{R}(x,w^{(i)})}\cdot\mathrm{Ext}_{R}(x,w^{(m)}).$ (85)
In the standard self-reducibility framework, we estimate each ratio $\rho_{i}=\mathrm{Ext}_{R}(x,w^{(i-1)})/\mathrm{Ext}_{R}(x,w^{(i)})$ using the assumed approximation scheme. Let $A_{i}$ be the estimator for the $i$-th ratio such that:
$\mathbb{P}\left[\frac{1}{r}\leq\frac{A_{i}}{\rho_{i}}\leq r\right]\geq 1-\frac{\delta}{m}.$ (86)
By the union bound, the probability that at least one of these $m$ estimations fails to stay within the ratio $r$ is at most $m\cdot(\delta/m)=\delta$. Therefore, with probability at least $1-\delta$, all $m$ estimations are successful. Under this condition, the final estimate $\hat{N}=\prod_{i=1}^{m}A_{i}$ satisfies:
$\frac{1}{r^{m}}\leq\frac{\hat{N}}{N_{R}(x)}\leq r^{m}.$ (87)
From Proposition C.14, we have $r^{m}=(1+\frac{\varepsilon}{2m})^{m}<1+\varepsilon$. Furthermore, for $0<\varepsilon\leq 1$, it holds that $1/r^{m}>(1+\varepsilon)^{-1}\geq 1-\varepsilon$, ensuring a relative error of at most $\varepsilon$.
Regarding complexity, each of the $m$ calls to the FPRAS for $\mathrm{Ext}_{R}$ runs in time $\mathrm{poly}(n,(r-1)^{-1},\log(m/\delta))$. Substituting $r-1=\varepsilon/2m$, the runtime per call is $\mathrm{poly}(n,m/\varepsilon,\log(m/\delta))$. Since $m=p(n)$ is a polynomial in $n$, the total running time is $\mathrm{poly}(n,1/\varepsilon,\log(1/\delta))$, which satisfies the requirements for an FPRAS. ∎
We have shown that approximate counting reduces to approximating the extension counting function. For a general relation $R$, however, the connection between the extension counting function and the original counting problem for $R$ remains unclear. This gap is bridged by the notion of self-reducibility: for self-reducible relations, the extension counting function can be reduced back to the original counting problem for $R$.
Definition C.15 (Schnorr, 1976).
@@
For self-reducible relations, the extension counting function is no harder to approximate than the original counting problem.
Proposition C.16 (Jerrum et al., 1986).
Let $R$ be self-reducible. If there exists an FPRAS for $N_{R}$, then there exists an FPRAS for $\mathrm{Ext}_{R}$.
Proof.
By the definition of self-reducibility, the extension function $\mathrm{Ext}_{R}(x,w)$, which counts the number of strings $y$ such that $(x,wy)\in R$, can be mapped to the counting problem of a modified instance. Specifically, for any prefix $w$ where $|w|\leq\sigma(x)$, there exists a polynomial-time computable mapping $\psi$ such that:
$\mathrm{Ext}_{R}(x,w)=|\{y\in\Sigma^{\sigma(x)-|w|}:(x,wy)\in R\}|=N_{R}(\psi(x,w)).$ (92)
Since $R$ is self-reducible, the instance $x_{w}=\psi(x,w)$ can be constructed in polynomial time relative to $|x|$. By the hypothesis, there exists an FPRAS for $N_{R}$, which provides a randomized $(1\pm\epsilon)$-approximation of $N_{R}(x_{w})$ in time polynomial in $|x_{w}|$ and $1/\epsilon$. Consequently, this algorithm serves as an FPRAS for $\mathrm{Ext}_{R}(x,w)$, as it runs in polynomial time and satisfies the required approximation guarantees. ∎
C.4 Proof for Theorem 4.3
Lemma C.17 (Formal Statement of Lemma 4.3).
Assume that $\mathsf{FPTAS}\subsetneq\mathsf{FPRAS}$ for self-reducible relations. There exists a self-reducible relation $R$ and an associated function $\mathrm{Ext}_{R}:\Sigma^{*}\times\Sigma^{*}\to\mathbb{N}$ defined by $\mathrm{Ext}_{R}(x,y_{<i})\coloneqq|\{\,z\in\Sigma^{*}:(x,y_{<i}z)\in R\,\}|$ such that language models with CoT using polynomially many reasoning steps, which output a distribution for a given input $(x,y_{<i})\in\Sigma^{n}\times\Sigma^{*}$ by using a linear head for the last hidden state before emitting the output token, admit an FPRAS for $\mathrm{Ext}_{R}$, whereas no latent thought with polynomially many iterations admits the same approximation guarantee using a linear head for the last hidden state before emitting the output token.
Proof.
By Lemma C.6, CoT can simulate any probabilistic Turing machine running in polynomial time; thus, it can implement an FPRAS for $N_{R}$. By Proposition C.16, the existence of an FPRAS for $N_{R}$ further implies the existence of an FPRAS for the extension function $\mathrm{Ext}_{R}$. On the other hand, latent thought consisting of a polynomial number of iterations can always be simulated by a deterministic polynomial-time Turing machine, provided that all state transitions in the latent computation are deterministic. Consequently, if such a latent thought process were to admit an FPRAS for $\mathrm{Ext}_{R}$, it would effectively yield a deterministic polynomial-time approximation scheme. By Proposition C.16, the existence of such a scheme for $\mathrm{Ext}_{R}$ would imply the existence of an FPTAS for $N_{R}$. However, under the standard complexity-theoretic assumption that $\mathsf{FPTAS}\subsetneq\mathsf{FPRAS}$ for self-reducible relations, there exists a self-reducible relation $R$ whose counting function admits an FPRAS but no FPTAS. This yields a contradiction. ∎
C.5 Proof for Theorem 4.4
Definition C.18 (FPAUS).
Uniform generation asks to sample an element $y$ uniformly at random from $R(x)$. A fully polynomial almost uniform sampler (FPAUS) for $R$ is a randomized algorithm that, given an input $x\in\Sigma^{*}$ and an accuracy parameter $\varepsilon>0$, runs in time polynomial in $|x|$ and $\log(1/\varepsilon)$, and outputs a distribution $q(\cdot\mid x)$ such that
where $U(R(x))$ denotes the uniform distribution over the set $R(x)$, and $\|\cdot\|_{\mathrm{TV}}$ denotes total variation distance.
For self-reducible relations, the following holds.
Theorem C.19 (Jerrum et al., 1986).
@@
Proof of Theorem 4.4.
The target uniform conditional distribution is defined as follows:
@@
Assume an FPRAS $\mathcal{A}(x,\varepsilon,\delta)$ exists for the self-reducible relation $|R(x)|$. By Proposition C.16, the existence of an FPRAS for $|R(x)|$ implies the existence of an FPRAS for the extension function $\operatorname{EXT}_{R}$. We construct a CoT that samples $y\in R(x)$ by sequentially approximating these conditional probabilities. For each step $i\in\{1,\dots,m(n)\}$, the CoT computes $(1\pm\varepsilon)$-accurate estimates $\widehat{\operatorname{EXT}}_{R}(x,y_{<i}u)$ for all $u\in\Sigma$ using $\mathcal{A}$, and induces the following distribution:
$\pi(y_{i}\mid x,y_{<i})\;:=\;\frac{\widehat{\operatorname{EXT}}_{R}(x,y_{<i}y_{i})}{\sum_{u\in\Sigma}\widehat{\operatorname{EXT}}_{R}(x,y_{<i}u)}.$ (94)